T ELEIRO ILU, U UNUSTAPLE 1. An entrepreneur must have $25,000 in 3 years to start a new business. (a) Calculate the amount tplease answer correctly thanks
Answer #1

rate positively ..

Ans 1 a)
Amount to deposit = 25000/(1+6%)^3
20990.482
Ans 1 b) Financial function =
=PV(6%,3,0,-25000,0) $20,990.48
Ans 2 a)
Year Payment PVIF @ 9% present value
1 7000          0.9174      6,422.02
2 7000          0.8417      5,891.76
3 7000          0.7722      5,405.28
4 7000          0.7084      4,958.98
5 7000          0.6499      4,549.52
6 7000          0.5963      4,173.87
7 7000          0.5470      3,829.24
8 7000          0.5019      3,513.06
   38,743.73
Ans = 38,743.73
Ans 2 b) Financial function =
=PV(9%,8,-7000,0,0) $38,743.73